OpenStudy (nikki17_guess):

Bank A charges a monthly service fee of $5.00 and a per-check fee of $0.15, while Bank B charges a monthly service fee of $2.50 and a per-check fee of $0.35. Which of these numbers of monthly checks will cause Bank A to charge less in fees than Bank B? A.5 B.19 C.10 D.11

OpenStudy (photon336):

\[Bank~A~y= 5.00+0.15c\] \[Bank~B~y = 2.50+0.35c\]

OpenStudy (nikki17_guess):

i don't get it

OpenStudy (photon336):

Then what we can do is set up an inequality

OpenStudy (photon336):

Bank A < Bank B 5.00+0.15c < 2.50+0.35c 2.50+0.15c < 0.35c -0.15c -0.15c 2.50 < 0.20c 12.5 < c or c > 12.5

OpenStudy (photon336):

so for bank a to charge a lower fee than bank B you need to have deposited more than 13 checks.

OpenStudy (photon336):

@nikki17_guess

OpenStudy (nikki17_guess):

oh ok thank you cx

OpenStudy (photon336):

no problem